Butane (C4H10) has a heat of combustion of -2882 kJ/mol. If a 10.09 g sample of butane is burned in a bomb calorimeter containing 2.610 kg of water initially at 20.13 degrees C, find the final water temperature.
Answer #1

58 g butane gives 2882 KJ energy  

10.09 g butane gives 10.09 x 2882 / 58 = 501.37 KJ energy

now this energy used to raise the temperature of water

q = mC\DeltaT

501370 J = 2610 g x 4.184 J/g0C x \Delta T

\DeltaT = 501370 / 10920

\DeltaT = 45.9

Tf = 21.13 + 45.9 = 66.03

answer = 66.03 0C